If ‘S’ is stress and ‘Y’ is young’s modulus of material of a wire, the energy stored in the wire per unit volume is
(a) S2/2Y
(b) 2S2Y
(c) S/2Y
(d) 2Y/S2

Grade:12th pass

3 Answers

Aditi Chauhan
9 years ago

(a) Energy stored per unit volume,

E = 1/2 * stress * strain

We know that,

E = 1/2 * stress * stress/Y = 1/2 .S2/Y

ginny
5 years ago
y= stress /strain
strain=stress/y
e= ½ *stress*strain
e= 1/2(stress*stress)/y
e=1/2 s^2/y
the answer is option (a)
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

y= stress /strain
strain=stress/y
e = ½ *stress*strain
e = 1/2(stress*stress)/y
e = 1/2 s^2/y
e = s^2/2y

Thanks and Regards